DVIBOOK(1) DVIBOOK(1) NAME dvibook−rearrangepagesinDVIfileintosignatures SYNOPSIS dvibook[−q][−s<num>][−iinfile][−ooutfile][infile[outfile]] DESCRIPTION DvibookrearrangespagesfromaDVIfileproducedbyTeXinto‘‘signatures’’forprintingbooksor booklets,creatinganewDVIfileusablebyanyoftheTeXconversionprograms. The−soptionselectsthesizeofsignaturewhichwillbeused.Thesignaturesizeisthenumberof sideswhichwillbefoldedandboundtogether;thenumbergivenshouldbeamultipleoffour.The defaultistouseonesignatureforthewholefile.Extrablanksideswillbeaddedifthefiledoes notcontainamultipleoffourpages. Dvibooknormallyprintsthepagenumbersofthepagesrearranged;the−qoptionsuppressesthis. AUTHOR AngusDuggan,fromdviselect(1),byChrisTorek,UniversityofMaryland SEEALSO dviconcat(1),dviselect(1),latex(1),tex(1) MC-TeXUser’sGuide TheTeXbookDvibookdoesnotadjusttheparametersinthepostamble;however,sincethesevaluesarenormallyusedonlytosizecertainstructuresintheoutputconversionprograms,andthe parametersneverneedtobeadjustedupward,thishasnotproventobeaproblem.